Directions:Write the equation of the line
Question: has slope 2/3 and passes through (1,3)
My answer: y-3=2/3(x-1)
Is that correct?

Answer by Boreal(15235)

you can substitute in 1 for x and 3 for y and see if it works
3-3=(2/3)(0), which is true but that isn't the equation of the line in slope intercept form
point slope formula y-y1=m(x-x1), m slope, and (x1, y1) is point
y-3=(2/3)(x-1) which is what you have.
that is in point-slope form
y=(2/3)x-2/3+3
y=(2/3)x+(7/3) is slope-intercept form. It depends upon what is being asked for.